OpenStudy (anonymous):

Solve. x divided by eight equals five divided by forty x equals one x equals five-eighths x equals one and three-fifths x equals five

OpenStudy (anonymous):

its just an excuse to b on open study

OpenStudy (anonymous):

x=1, so the first option

OpenStudy (texaschic101):

x/8 = 5/40 cross multiply 40x = 40 -- divide by 40 x = 1

OpenStudy (anonymous):

i know what to do read the first thing i put

OpenStudy (texaschic101):

you really need an excuse to be on here ?

OpenStudy (anonymous):

kinda